OpenStudy (anonymous):

What is the sum of the angle measures in a polygon that has 20 sides?

OpenStudy (saifoo.khan):

use the formula.

OpenStudy (saifoo.khan):

\[(x-2)*180\]

OpenStudy (saifoo.khan):

where x is the number of sides.

OpenStudy (saifoo.khan):

are the number of sides*

OpenStudy (anonymous):

180(x-2) 180(18) 3240degrees

OpenStudy (saifoo.khan):

Runner, no it isnt. LOL.

OpenStudy (saifoo.khan):

its for 4 sided.

OpenStudy (anonymous):

Sum of exterior angles of any polygon is 360 degrees. Each exterior angle of a regular polygon = 360/n where n is the number of sides. So, n = 360/20 = 18. OR Each interior angle of the regular polygon is 180-20 = 160 degrees So, sum of interior angles of the regular polygon = 160n degrees where n is the number of angles (also the number of sides) Now, sum of interior angles of any polygon = (n-2)x180 degrees Thus 160n = (n-2)x180 or 160n = 180n - 360 Solving for n, we get n = 18.